OpenStudy (brittany16):

Which operation would isolate the variable in this equation? 5/6m=-30

OpenStudy (anonymous):

multiply both sides by the reciprocal of 5/6, 6/5 m=

OpenStudy (brittany16):

m-what?

OpenStudy (anonymous):

m=-36

OpenStudy (anonymous):

5/6m(6/5)=-30(6/5) m=36

jhonyy9 (jhonyy9):

multiply both sides by 6m and get 5= (-30)*6m 5= -180m 5/180=-m 36=-m -36=m